# A chaotic lattice field theory in two dimensions

### v2

@@ XI Shadowing
It was shown numerically by Gutkin et al. [36, 37] that pairs of interior alphabet Eq. (165) spatiotemporal cat periodic states of a fixed spatial width $L$ that share sets of sub-mosaics, shadow each other when evolved forward-in-time. Here, in Sec. XI.2, we check numerically spatiotemporal cat shadowing for arbitrary periodic states, without alphabet restrictions, and without any time evolution. Intuitively, if two unstable periodic states $\mathsf{\Phi}$, $\mathsf{\Phi}^{\prime}$ share a common sub-mosaic $\mathsf{M}$ of volume ${N}_{\mathsf{M}}$, they shadow each other with exponential accuracy of order of $\propto\exp(-\lambda{N}_{\mathsf{M}})$. In time-evolution formulation, $\lambda$ is the leading Lyapunov exponent. What is it for spatiotemporal systems?
@@
XI.1 Shadowing, one-dimensional temporal cat
@@
XI.2 Shadowing, two-dimensional spatiotemporal cat
@@
The two-dimensional $\mu^{2}=1$ spatiotemporal cat Eq. (49) periodic states are labelled by two-dimensional mosaics, 8-letter alphabet Eq. (163), as in Fig. 12.
@@
(a)    Refer to caption
(b)    Refer to caption
Figure 12: (Color online) Mosaics Eq. (37) of two $18\!\times\!18]_{0}$ spatiotemporal cat periodic states which share the sub-mosaic within the $12\!\times\!12]$ region enclosed by the black square, and have different symbols outside the sub-mosaic. Color coded 8-letter alphabet Eq. (163), $\mu^{2}=1$. Continued in Fig. 13.
@@
Figure 13: (Color online) $\mu^{2}=1$ spatiotemporal cat. (a) The log of mean of point-wise field value distances $|\phi_{z}-\phi^{\prime}_{z}|$ over all lattice sites of $z\in[18\!\times\!18]$ primitive cell, averaged over the 250 pairs of periodic states, like the pair of Fig. 12. (b) The log of mean point-wise distances $|\phi_{9,t}-\phi^{\prime}_{9,t}|$ evaluated across the strip $z=(9,t)$, $t=1,2,\dots,18$, going through the center of the primitive cell. The decrease from edge to the center is approximately linear, with slope $\approx-1.079$.
XI.3 Green’s function of two-dimensional spatiotemporal cat
Mosaic $\mathsf{M}$ is admissible (see Sec. II.3) if field configuration $\mathsf{\Phi}_{\mathsf{M}}$ is a periodic state, i.e., all lattice site fields are confined to Eq. (162), the compact boson hypercube state space $\phi_{z}\in[0,1)$.
@@ XI.3 Green’s function of two-dimensional spatiotemporal cat
where $K_{0}$ is the modified Bessel function of the second kind. For large spacetime separations, $|x-x^{\prime}|\to\infty$, the asymptotic form of the Green’s function is
@@
In the numerical example of Sec. XI.2, we have set Klein-Gordon mass $\mu=1$, so the Green’s function of the continuum screened Poisson equation is a good approximation to the discrete spatiotemporal cat Green’s function, where the rate of decrease of correlations computed from the Fig. 13 (b) is approximately $\exp(-\mu^{\prime}r)$, with $\mu^{\prime}=-1.079$ the slope computed from the log plot of the mean distances of field values between shadowing periodic states.
XI.4 Convergence of evaluations of observables
Computed on primitive cells $\mathbb{A}$ of increasing volume ${N}_{\mathbb{A}}$, the expectation value of an observable (Sec. I.4) converges towards the exact, infinite Bravais lattice value (Sec. IX). As the simplest case of such sequence of primitive cell approximations, take a rectangular primitive cell ${L}\!\times\!{T}]_{0}$}, and evaluate stability exponents $\langle{\lambda}\rangle_{{[r{L}\!\times\!r{T}]}}$ (Sec. VIII.2) for the sequence of primitive cell repeats $r{L}\!\times\!r{T}]_{0}$} of increasing $r$.
That the convergence of such series of primitive cell approximations is a shadowing calculation can be seen by inspection of Fig. 11. The exact stability exponent $\lambda$} is obtained by integration over the bands (smooth surfaces in the figures). A shadowing approximation ${\lambda}_{{[{L}\!\times\!{T}]_{{S}}}}$ is a finite sum over primitive cells ${L}\!\times\!{T}]_{{S}}$}, black dots in the figures, that shadows the curved surface, with increasing accuracy as the primitive cell volume ${N}_{\mathbb{A}}$ increases. Here shadowing errors are Hipparchus’ errors Eq. (91) of replacing arcs by cords, as in approximating $2\pi$ by the perimeter of a regular $n$-gon. The sense in which such shadowing or ‘curvature’ errors are exponentially small for one-dimensional, temporal lattice chaotic systems is explained in Refs. 49, 182, 183. We have not extended such error estimates to the spatiotemporal case, so here we only present numerical evidence that they are exponentially small.
@@ XI.4 Convergence of evaluations of observables
For various primitive cell sequences of rectangular shapes ${L}\!\times\!{T}]_{0}$}, the stability exponents of repeat primitive cells $r{L}\!\times\!r{T}]_{0}$} also converge to $\lambda$} exponentially, with the same convergence rate $\approx 1.055\cdots$. We have no theoretical estimate of this rate, but it appears to be close to the Klein-Gordon mass $\mu=1$, within the shadowing error estimates of Sec. XI.2.
XII Summary and open questions
@@
These questions have been studied extensively for systems of small spatial extent, whose inertial manifold dimension is relatively small. [186, 187, 188, 189, 190] Recent experimental and theoretical advances [11, 15, 191, 14, 192] support such forward-in-time dynamical vision also for spatially extended, turbulent systems (see Ref. 193 for a gentle introduction). By now thousands of such ‘exact coherent structures’ have been computed, always confined to small spatial domains, while the flows of interest (pipe, channel, plane flows) are flows on infinite spatial domains. Going from spatially small to spatially infinite systems requires a radical shift in the point of view. To describe those, we have to recast equations such as the Navier-Stokes as a spacetime theory, with all infinite translational symmetry directions treated on equal footing. It is a bold leap, a theory of turbulence that does away with dynamics. But we have no choice. For spatially extended systems evolution forward in time is insanely unstable. [13, 194] Not only have time-evolution numerical codes [195, 196] not worked on large domains, in retrospect it is clear that they never could have worked.
@@
The first problem that we face is global: determining and organizing infinities of unstable multi-periodic states over $\infty$-dimensional state spaces, orbits that are presumed to form the skeleton of turbulence. We characterize and classify them by their shapes, captured by corresponding ‘mosaics’ (Sec. II.3). The feel is of statistical-mechanics, like enumeration of Ising configurations. For temporal evolution, in the large volume ${T}\to\infty$ limit our partition sum $Z$, Eq. (116), is Ruelle’s ‘partition function’, [19] our functional $W$, Eq. (29), is Ruelle’s ‘pressure’, [197] and our deterministic zeta function $\zeta$, Eq. (131), is Ruelle’s ‘dynamical zeta function’ [149] (for nomenclature, see ChaosBook  remark 20.2 ‘Pressure’). In the spatiotemporal theory, in the large volume ${L}{T}\to\infty$ limit our functional $W$ is Grassberger’s ‘density of metric entropy’ [4] and Politi, Torcini & Lepri’s chronotopic ‘entropy potential function’. [5, 6, 7] Furthermore, as had previously been done implicitly in cycle expansion calculations of observables, [198, 199, 200] see Sec. X.6, Politi et al. relate expectation values of observables by Legendre transforms of $W$ to appropriate ‘effective actions’ or ‘Gibbs free energies’.
@@ XII.1 What is new
The essential innovation of our approach is the computation of spatiotemporal stability exponents over infinite Bravais lattices, see Fig. 8, 9 and 11, rather than forward-in-time Floquet/Lyapunov stability of finite time solutions over rectangular spacetime primitive cells. We work in infinite spacetime, not a mix of modest or large spatial extent with either long time ergodic estimates, or compact periodic time solutions. The orbit Jacobian operator Eq. (69), Eq. (55), the orbit stability exponent $\lambda_{p}$ Eq. (111), and the surprisingly simple exact deterministic spatiotemporal zeta function Eq. (129) are defined democratically over spacetime lattice momenta (Pikovsky’s Bloch quasi-momenta and Bountis’ Floquet quasi-energies).
Historically, in all of the previous work on spatially extended systems, the time and space instabilities were treated asymmetrically. While the spatial stability was parametrized by spatial Bloch wave number $k_{1}$, the time (in)stability was either estimated by forward evolution-in-time numerical Lyapunov-Bloch exponents [25] $\lambda(k_{1})$ (the Oseledets multiplicative ergodic theorem), or by means of compact, finite primitive cell (Sec. VIII) periodic time solutions. Examples of the latter are spacetime periodic solutions of Kuramoto-Sivashinsky and Navier-Stokes PDEs studied in Ref. 211, 12, 187, 212, 213, 194, 214.
@@
And, finally: now that we have a hierarchy of multi-periodic states, what is the expectation value of any observable of the theory? The answer is given by deterministic partition sums and zeta functions of Sec. X. All chaoticity is due to the intrinsic instability of deterministic solutions, and our spatiotemporal deterministic partition sums Eq. (116) and zeta functions Eq. (129) are exact, not merely saddle points approximations to a theory. The issues in their applications are numerical: how many periodic states, evaluated to what accuracy have to be included into a truncated zeta function, to attain a desired accuracy?
Our deterministic field theory replaces numerical extrapolations of finite spacetime simulations of the 1990’s chronotopy by the exact deterministic zeta function, with computable cycle expansion truncations errors, decreasing exponentially with the spatiotemporal volumes of periodic states included in its evaluation. Our spatiotemporal deterministic zeta function, as far as we know, is new.
XII.2 Open questions
@@
    For pedagogical reasons –in order to start out with determinants of finite matrices, rather than immediately grapple with functional determinants– we have presented here the spacetime deterministic field theory in its discretized, $d$-dimensional lattice form. We expect the periodic orbit formulation of continuum spacetime theory to be of essentially of the same form, with the spacetime periodicities (Bravais lattice primitive vectors) defined over the continuum, $\mathbf{a}_{j}\in\mathbb{R}^{d}$, the stability exponent Eq. (111) evaluated as functional integral over its Brillouin zone, generating function variable $z$ replaced by a Laplace transform variable $s$, $z=e^{-s}$ (see ChaosBook  Eq. (21.20)), and the deterministic partition sum Eq. (116) of form
@@
    In Sec. V we have assumed that the only symmetry of the theory is the translation group $T$. However, one needs to quotient all spacetime and internal symmetries. For a one-dimensional lattice field theory we have done this in companion paper I, [16] and derived the dihedral-space group $G=\textrm{D}_{\infty}$ zeta function for time-reversal invariant field theories, drawing inspiration from Lind’s topological (rather than our weighted, ‘dynamical’) zeta function, [156]
@@
    a generalization of Artin-Mazur zeta function. In present context, $G$ is the crystallographic space group of a field theory over a hypercubic lattice $\mathbb{Z}^{d}$, $H$ a finite-index |G/H||G/{H}| subgroup of $G$, and ${N}_{H}$ is the number of the periodic states that are invariant under actions of the subgroup $H$.
@@
    Describe the admissible mosaics of a nonlinear field theory with pruning, i.e., with couplings weaker than those topologically equivalent to the anti-integrable limit (see Sec. II.3 and companion paper III [17]).
@@ Acknowledgements.
This paper has been profoundly influenced by anti-integrability viewpoint of Serge Aubry, whose 80th birthday it celebrates. The work of H. L. was fully supported, and of P. C. was in part supported by the family of late G. Robinson, Jr.. This research was initiated during the KITP UC Santa Barbara 2017 Recurrent Flows: The Clockwork Behind Turbulence program, supported in part by the National Science Foundation under Grant No. NSF PHY-1748958. We are grateful to S.D.V. Williams, M.N. Gudorf and X. Wang for many inspiring discussions. No actual cats, graduate or undergraduate, have shown interest in, or were harmed during this research.
@@ A.1 Examples of prime orbits
over primitive cell $2\!\times\!1]_{1}$}. As shown in Fig. 3 (a), Bravais lattice $2\!\times\!2]_{0}$ is a sublattice of $2\!\times\!1]_{1}$. Over the infinite spacetime $\mathsf{\Phi}$ and $\mathsf{\Phi}_{p}$ are the same field configuration, as is clear by inspection of Fig. 6 (c).

### v3

@@ XII Shadowing
It was shown numerically by Gutkin et al. [36, 37] that pairs of interior alphabet Eq. (165) spatiotemporal cat periodic states of a fixed spatial width $L$ that share sets of sub-mosaics, shadow each other when evolved forward-in-time. Here, in Sec. XII.2, we check numerically spatiotemporal cat shadowing for arbitrary periodic states, without alphabet restrictions, and without any time evolution. Intuitively, if two unstable periodic states $\mathsf{\Phi}$, $\mathsf{\Phi}^{\prime}$ share a common sub-mosaic $\mathsf{M}$ of volume ${N}_{\mathsf{M}}$, they shadow each other with exponential accuracy of order of $\propto\exp(-\lambda{N}_{\mathsf{M}})$. In time-evolution formulation, $\lambda$ is the leading Lyapunov exponent. What is it for spatiotemporal systems?
@@
XII.1 Shadowing, one-dimensional temporal cat
@@
XII.2 Shadowing, two-dimensional spatiotemporal cat
@@
The two-dimensional $\mu^{2}=1$ spatiotemporal cat Eq. (49) periodic states are labeled by two-dimensional mosaics, 8-letter alphabet Eq. (163), as in Fig. 5.
@@
Figure 13: $\mu^{2}=1$ spatiotemporal cat. (a) The log of mean of point-wise field value distances $|\phi_{z}-\phi^{\prime}_{z}|$ over all lattice sites of $z\in[18\!\times\!18]$ primitive cell, averaged over the 250 pairs of periodic states, like the pair of Fig. 5. (b) The log of mean point-wise distances $|\phi_{9,t}-\phi^{\prime}_{9,t}|$ evaluated across the strip $z=(9,t)$, $t=1,2,\dots,18$, going through the center of the primitive cell. The decrease from edge to the center is approximately linear, with slope $\approx-1.079$.
XII.3 Green’s function of two-dimensional spatiotemporal cat
Mosaic $\mathsf{M}$ is admissible (see Sec. III.3) if field configuration $\mathsf{\Phi}_{\mathsf{M}}$ is a periodic state, i.e., all lattice site fields are confined to Eq. (162), the compact boson hypercube state space $\phi_{z}\in[0,1)$.
@@ XII.3 Green’s function of two-dimensional spatiotemporal cat
where $K_{0}$ is the modified Bessel function of the second kind. [193] For large spacetime separations, $|x-x^{\prime}|\to\infty$, the asymptotic form of the Green’s function is
@@
In the numerical example of Sec. XII.2, we have set Klein-Gordon mass $\mu=1$, so the Green’s function of the continuum screened Poisson equation is a good approximation to the discrete spatiotemporal cat Green’s function, where the rate of decrease of correlations computed from the Fig. 13 (b) is approximately $\exp(-\mu^{\prime}r)$, with $\mu^{\prime}=-1.079$ the slope computed from the log plot of the mean distances of field values between shadowing periodic states.
XII.4 Convergence of evaluations of observables
Computed on primitive cells $\mathbb{A}$ of increasing volume ${N}_{\mathbb{A}}$, the expectation value of an observable (Sec. II.4) converges towards the exact, infinite Bravais lattice value (Sec. X). As the simplest case of such sequence of primitive cell approximations, take a rectangular primitive cell ${L}\!\times\!{T}]_{0}$}, and evaluate stability exponents $\langle{\lambda}\rangle_{{[r{L}\!\times\!r{T}]}}$ (Sec. IX.2) for the sequence of primitive cell repeats $r{L}\!\times\!r{T}]_{0}$} of increasing $r$.
That the convergence of such series of primitive cell approximations is a shadowing calculation can be seen by inspection of Fig. 10. The exact stability exponent $\lambda$} is obtained by integration over the bands (smooth surfaces in the figures). A shadowing approximation ${\lambda}_{{[{L}\!\times\!{T}]_{{S}}}}$ is a finite sum over primitive cells ${L}\!\times\!{T}]_{{S}}$}, black dots in the figures, that shadows the curved surface, with increasing accuracy as the primitive cell volume ${N}_{\mathbb{A}}$ increases. Here shadowing errors are Hipparchus’ errors Eq. (91) of replacing arcs by cords, as in approximating $2\pi$ by the perimeter of a regular $n$-gon. The sense in which such shadowing or ‘curvature’ errors are exponentially small for one-dimensional, temporal lattice chaotic systems is explained in Refs. 49, 194, 195. We have not extended such error estimates to the spatiotemporal case, so here we only present numerical evidence that they are exponentially small.
@@ XII.4 Convergence of evaluations of observables
For various primitive cell sequences of rectangular shapes ${L}\!\times\!{T}]_{0}$}, the stability exponents of repeat primitive cells $r{L}\!\times\!r{T}]_{0}$} also converge to $\lambda$} exponentially, with the same convergence rate $\approx 1.055\cdots$. We have no theoretical estimate of this rate, but it appears to be close to the Klein-Gordon mass $\mu=1$, within the shadowing error estimates of Sec. XII.2.
XIII Summary and open questions
@@
These questions have been studied extensively for systems of small spatial extent, whose inertial manifold dimension is relatively small. [198, 199, 200, 201, 202] Recent experimental and theoretical advances [11, 15, 203, 14, 204] support such forward-in-time dynamical vision also for spatially extended, turbulent systems (see Ref. 205 for a gentle introduction). By now thousands of such ‘exact coherent structures’ have been computed, always confined to small spatial domains, while the flows of interest (pipe, channel, plane flows) are flows on infinite spatial domains. Going from spatially small to spatially infinite systems requires a radical shift in the point of view. To describe those, we have to recast equations such as the Navier-Stokes as a spacetime theory, with all infinite translational symmetry directions treated on equal footing. It is a bold leap, a theory of turbulence that does away with dynamics. However, we have no choice. For spatially extended systems evolution forward in time is insanely unstable. [13, 206] Not only have time-evolution numerical codes [207, 208] not worked on large domains, in retrospect it is clear that they never could have worked.
@@
The first problem that we face is global: determining and organizing infinities of unstable multi-periodic states over $\infty$-dimensional state spaces, orbits that are presumed to form the skeleton of turbulence. We characterize and classify them by their shapes, captured by corresponding ‘mosaics’ (Sec. III.3). The feel is of statistical-mechanics, such as enumeration of Ising configurations. For temporal evolution, in the large volume ${T}\to\infty$ limit, our partition sum $Z$, Eq. (116), is Ruelle’s ‘partition function’, [19] our functional $W$, Eq. (29), is Ruelle’s ‘pressure’, [209] and our deterministic zeta function $\zeta$, Eq. (131), is Ruelle’s ‘dynamical zeta function’ [155] (for nomenclature, see remark 20.2 ‘Pressure’ in Ref. 210). In the spatiotemporal theory, in the large volume ${L}{T}\to\infty$ limit our functional $W$ is Grassberger’s ‘density of metric entropy’ [4] and Politi, Torcini & Lepri’s chronotopic ‘entropy potential function’. [5, 6, 7] Furthermore, as had previously been done implicitly in cycle expansion calculations of observables, [211, 212, 213] see Sec. XI.6, Politi et al. relate expectation values of observables by Legendre transforms of $W$ to appropriate ‘effective actions’ or ‘Gibbs free energies’.
@@ XIII.1 What is new
The essential innovation of our approach is the computation of spatiotemporal stability exponents over infinite Bravais lattices, see Fig. 1, 11 and 10, rather than forward-in-time Floquet/Lyapunov stability of finite time solutions over rectangular spacetime primitive cells. We work in infinite spacetime, not a mix of modest or large spatial extent with either long time ergodic estimates, or compact periodic time solutions. The orbit Jacobian operator Eq. (69), Eq. (55), the orbit stability exponent $\lambda_{p}$ Eq. (111), and the surprisingly simple exact deterministic spatiotemporal zeta function Eq. (129) are defined democratically over spacetime lattice momenta (Pikovsky’s Bloch quasi-momenta and Bountis’ Floquet quasi-energies).
Historically, in all of the previous work on spatially extended systems, the time and space instabilities were treated asymmetrically. While the spatial stability was parameterized by spatial Bloch wave number $k_{1}$, the time (in)stability was either estimated by forward evolution-in-time numerical Lyapunov-Bloch exponents [25] $\lambda(k_{1})$ (the Oseledets multiplicative ergodic theorem), or by means of compact, finite primitive cell (Sec. IX) periodic time solutions. Examples of the latter are spacetime periodic solutions of Kuramoto-Sivashinsky and Navier-Stokes PDEs studied in Ref. 224, 12, 199, 225, 226, 206, 227.
@@
And, finally: now that we have a hierarchy of multi-periodic states, what is the expectation value of any observable of the theory? The answer is given by deterministic partition sums and zeta functions of Sec. XI. All chaoticity is due to the intrinsic instability of deterministic solutions, and our spatiotemporal deterministic partition sums Eq. (116) and zeta functions Eq. (129) are exact, not merely saddle point approximations to a theory. The issues in their applications are numerical: how many periodic states, evaluated to what accuracy have to be included into a truncated zeta function, to attain a desired accuracy?
Our deterministic field theory replaces numerical extrapolations of finite spacetime simulations of the 1990s chronotopy by the exact deterministic zeta function, with computable cycle expansion truncations errors, decreasing exponentially with the spatiotemporal volumes of periodic states included in its evaluation. Our spatiotemporal deterministic zeta function, as far as we know, is new.
XIII.2 Open questions
@@
    For pedagogical reasons –in order to start out with determinants of finite matrices, rather than immediately grapple with functional determinants– we have presented here the spacetime deterministic field theory in its discretized, $d$-dimensional lattice form. We expect the periodic orbit formulation of continuum spacetime theory to be of essentially of the same form, with the spacetime periodicities (Bravais lattice primitive vectors) defined over the continuum, $\mathbf{a}_{j}\in\mathbb{R}^{d}$, the stability exponent Eq. (111) evaluated as a functional integral over its Brillouin zone, generating function variable $z$ replaced by a Laplace transform variable $s$, $z=e^{-s}$ (see ChaosBook  Eq. (21.20) in Ref. 158), and the deterministic partition sum Eq. (116) of the form
@@
    In Sec. VI we have assumed that the only symmetry of the theory is the translation group $T$. However, one needs to quotient all spacetime and internal symmetries. For a one-dimensional lattice field theory we have done this in companion paper I, [16] and derived the dihedral space group $G=\textrm{D}_{\infty}$ zeta function for time-reversal invariant field theories, drawing inspiration from Lind’s topological (rather than our weighted, ‘dynamical’) zeta function, [165]
@@
    a generalization of Artin-Mazur zeta function. In the present context, $G$ is the crystallographic space group of a field theory over a hypercubic lattice $\mathbb{Z}^{d}$, $H$ a finite-index |G/H||G/{H}| subgroup of $G$, and ${N}_{H}$ is the number of the periodic states that are invariant under actions of the subgroup $H$.
@@
    Describe the admissible mosaics of a nonlinear field theory with pruning, i.e., with couplings weaker than those topologically equivalent to the anti-integrable limit (see Sec. III.3 and companion paper III [17]).
@@ Acknowledgements.
This paper has been profoundly influenced by anti-integrability viewpoint of Serge Aubry, whose 80th birthday it celebrates. The work of H. L. was fully supported, and of P. C. was in part supported by the family of late G. Robinson, Jr. This research was initiated during the KITP UC Santa Barbara 2017 Recurrent Flows: The Clockwork Behind Turbulence program, supported in part by the National Science Foundation under Grant No. NSF PHY-1748958. We are grateful to S.D.V. Williams, M.N. Gudorf and X. Wang for many inspiring discussions. No actual cats, graduate or undergraduate, have shown interest in, or were harmed during this research.
@@ A.1 Examples of prime orbits
over primitive cell $2\!\times\!1]_{1}$}. As shown in Fig. 6 (a), the Bravais lattice $2\!\times\!2]_{0}$ is a sublattice of $2\!\times\!1]_{1}$. Over the infinite spacetime $\mathsf{\Phi}$ and $\mathsf{\Phi}_{p}$ are the same field configuration, as is clear by inspection of Fig. 4 (c).